OpenStudy (thesecret20111):

1/a+1/v = 1/F Isolate for F showing steps. Please help!

OpenStudy (blurbendy):

F(a + v) = av F = av / (a +v)

OpenStudy (anonymous):

\[\frac{ 1 }{ a }+\frac{ 1 }{ v }=\frac{ 1 }{ F }\] \[F(\frac{1}{a}+\frac{1}{v})=1\] \[F(\frac{a+v}{av})=1\] \[\large F=\dfrac{ \dfrac{ 1 }{ a+v } }{ av }\] \[\large F=\dfrac{ av }{ a+v }\]

OpenStudy (thesecret20111):

Thanks so much guys :)